Please help me solve this question
Use Pythagorean theorem to find the length of the missing side. then find indicated trigonometric function of the given triangle . give a exact answer with a rational denominator
right triangle

theta of B

a=5
b=2
I started to do this question
c^2=5^2+2^2
c^2=25+4
c^2=29
c=Sqrt of 29
after that i am confused?

Answer by Alan3354(69443)

Use Pythagorean theorem to find the length of the missing side. then find indicated trigonometric function of the given triangle . give a exact answer with a rational denominator
right triangle

theta of B

a=5
b=2
I started to do this question
c^2=5^2+2^2
c^2=25+4
c^2=29
c=Sqrt of 29
after that i am confused?
==========
It's done
c = sqrt(29)
=====================
IDK what "theta of B" means.